OpenStudy (anonymous):

on wednesday olga's parents said she owed them too much money to borrow any more. on thursday she paid her parents $15 she had earned babysitting. on friday she borrowed $5 to go to a movie. on saturday she paid them the $12 she earned babysitting, then her debt was down $22. how much did she owe on wednesday? A. $22 B. $44 C. $52 D. $36

OpenStudy (ivettef365):

so you have the following x -15 +5 -12 = 22

OpenStudy (ivettef365):

solve for x

OpenStudy (anonymous):

Let amount owed on wednesday be x. Amount owed on thursday = x - 15 Amount owed on friday = x - 15 + 5 Amount owed on saturday = x - 15 + 5 - 12 = -22 Its helpful to let a certain unknown be x for simple forward calculation instead of reverse engineering.

OpenStudy (anonymous):

I got B.44 x-15+5-12=22 x-15+5=34 I add 12 x-15=29 I subtract 5 x=44 I add 15

OpenStudy (anonymous):

@melanie083194 did u get the answer all ready

OpenStudy (anonymous):

yea i got 44

OpenStudy (anonymous):

ok